Brown's, a local​ bakery, is worried about increased costs particularly energy. Last​ year's records can provide a fairly good estimate of the parameters for this year. Wende​ Brown, the​ owner, does not believe things have changed​ much, but she did invest an additional $2,500 for modifications to the​ bakery's ovens to make them more energy efficient. The modifications were supposed to make the ovens at least 15% more efficient. Brown has asked you to check the energy savings of the new ovens and also to look over other measures of the​ bakery's productivity to see if the modifications were beneficial. You have the following data to work​ with:

                                                                                             

Last Year

Now

Production​ (dozen)

1,500

1,500

Labor​ (hours)

340

325

Capital Investment​ ($)

15,000

17,500

Energy​ (BTU)

3,000

2,800

Energy productivity increase​ = ______%

​(enter your response as a percentage rounded to two decimal places and include a minus sign if​ necessary).

Captial productivity increase _____%

​(enter your response as a percentage rounded to two decimal places and include a minus sign if​ necessary).

labor productivity increase ______%

​(enter your response as a percentage rounded to two decimal places and include a minus sign if​ necessary).
Answer #1

Energy Prodictivity:

Last year: (1500*12) /3000 = 6.0 loaves/BTU

Now: (1500*12) /2800 = 6.43 loaves/BTU

Percent Change: ((6.43-6)/6) *100= 7.16%

Energy productivity increase​= 7.16%

There has definitely been an increase in the energy productivity which means there have been savings in energy.

Capital productivity:

Last year: 1500*12/15000 = 1.2 loaves/$ investment

Now: 1500*12/17500 = 1.03 loaves/$ investment

Percent Change: (1.03-1.2)/1.2*100= -14.17%

However as we can see there have been a decrease in the investment productivity.

Captial productivity increase -14.17%

Labor productivity:

Last year: 1500*12/340 = 52.94 loaves/labor hour

Now: 1500*12/325 = 55.38 loaves/labor hour

Percent Change: (55.38-52.94)/52.94*100= 4.60 %

There is also an increase in labor productivity

Labor productivity increase 4.60%